Tell Me Why Co. is expected to maintain a constant 4.21 percent growth rate in its dividends indefinitely. If the company has a dividend yield of 8.63 percent, what is the required return on the company’s stock?
Answer #1

9= 4.210% DI 8.63 l Required turn return = 1 +g = 8.63 + 4.21 = 12.84% Ans